# A Relative Homology Theory of Representation in Neural Networks

### v2

@@ C.1 Non-linear curves simulations
500 MSE 1e-4 1000 MSE $<$ 0.00002 50 3 1
C.2 Reproduction of the results of Naitzat et al.
As mentioned before, we reproduced the results from (Naitzat et al., 2020) with fewer samples due to the high computational demand of computing the overlap decomposition. This meant that we also had to reproduce the pre-processing steps in their pipeline. In order to speed up the computation of persistent homology and attain a more robust measure of distance, Naitzat et al. fixed a k-nearest neighbors graph at a fixed persistence parameter $\epsilon$ for all datasets. Since the homology groups of the datasets were known a priori, they were able to choose these values on the dataset and carry them forward in the computation of homology groups in deeper layers.
@@
D-I 7800/2200 CSE 2e-5 5000 Accuracy $>$ 0.999 30 15 9
D-II 7500/2500 CSE 2e-5 5000 Accuracy $>$ 0.999 30 15 9
D-III 8000/4000 CSE 2e-5 5000 Accuracy $>$ 0.999 30 15 9
In addition we wanted to avoid issues in homology estimation due to missclassified points, so we excluded all such points from the test set before performing a homology calculation. Given that our networks had almost perfect accuracy, this procedure removed very few points ($7.5\pm 12.9$). For the calculation of persistent homology we used the Ripser package (Tralie et al., 2018; Bauer, 2021).
C.3 Counting regions in the spheres dataset
@@ Appendix D Supplementary figures
Figure 5: More examples of non-linear manifolds and their homology groups. The numbers in the top right of each manifold show the ground truth number of holes. While our relative homology approach typically performs better, notice that in the fourth and (arguably) eighth example we see a homological type 2 error.

### v3

@@ C.2 Non-linear curves simulations
500 MSE 1e-4 1000 MSE << 0.00002 50 3 1
C.3 Reproduction of the results of Naitzat et al.
As mentioned before, we reproduced the results from Naitzat et al., (2020) with fewer samples due to the high computational demand of computing the overlap decomposition. This meant that we also had to reproduce the pre-processing steps in their pipeline. In order to speed up the computation of persistent homology and attain a more robust measure of distance, Naitzat et al. fixed a $k$-nearest neighbors graph at a fixed persistence parameter $\epsilon$ for all datasets. Since the homology groups of the datasets were known a priori, they were able to choose these values on the dataset and carry them forward in the computation of homology groups in deeper layers.
@@
D-I 7800/2200 CSE 2e-5 5000 Accuracy >> 0.999 30 15 9
D-II 7500/2500 CSE 2e-5 5000 Accuracy >> 0.999 30 15 9
D-III 8000/4000 CSE 2e-5 5000 Accuracy >> 0.999 30 15 9
In addition we wanted to avoid issues in homology estimation due to missclassified points, so we excluded all such points from the test set before performing a homology calculation. Given that our networks had almost perfect accuracy, this procedure removed very few points ($7.5\pm 12.9$). For the calculation of persistent homology we used the Ripser package (Tralie et al., (2018); Bauer, (2021)).
C.4 Counting regions in the spheres dataset
@@ Appendix D Supplementary figures
Figure 5: More examples of non-linear manifolds and their homology groups. The numbers in the top right of each manifold show the ground truth number of holes. While our quotient homology approach typically performs better, notice that in the fourth and (arguably) eighth example we see a homological type 2 error.
Figure 6: Histograms of the number of points that fall in each polyhedron induced by four layer networks trained and evaluated on all 60000 MNIST digits as a function of width. The colors (blue, orange, green and red) correspond to the respective four layers. One can observe that as the width of the network increases, we see an improvement in performance accompanied by fewer points belonging to each polyhedron. In layers three and four of the widest network, we see that each polyhedron is populated by only one point. This indicates that as networks become wider, the chance of only having convex intersections increases which is consistent with the fact that the volume of the average polyhedron decreases.
Figure 7: (Top left) Time in seconds to compute the H-representations of all polyhedra in a two layer network as function of width and input dimension on 1000 points sampled from the datasets in section 4.3. The remaining plots show the time it takes to compute the overlap decomposition as a function of width, input dimension and the sensitivity parameter. As can be seen the dependence on width is relatively minor. On the other hand, low values of the sensitivity parameter lead to very fast computing times, at the risk of missing important topological structure. All runs were performed on an Apple M3 Pro CPU with 11 cores.